OpenStudy (anonymous):

(9 – 8)2 × (24 – 6 3)

pooja195 (pooja195):

=(1)*2 * (-39) =2*(-39) =-78

OpenStudy (anonymous):

= (9 - 8)²(24 - 63) = 1²(- 39) = - 39